(1) What is the excess number of electrons that must be placed on each of two small sphere spaced 3cm apart so that a force of repulsion that acts between them becomes 10-19N


(a)25


(b)225


(c)625


(d)125


Please show the solution !!!!!!!!!!!!

Grade:11

4 Answers

Abhishekh kumar sharma
10 years ago

225

saurabh gupta
10 years ago

here q1=q2

10-19=(9*109)(q1.q2)/32

10-28=q1q2

10-14C=q1

 

q=ne

10-14=n*1.6*10-19

n=625

BAYANA SAGAR
10 years ago

I think that (c)is the correct answer.because no of electrons is 6.25( -18).10  

Mahendra Chapagain
5 years ago
625
solution:
we know F=K.qq/r^2
applying in above question F=K.(ne)^2/r^2
Now N^2=F.r^2/(e^2.K) where k is constant value 8.98*10^9
finally by solving we get n=625